What two numbers add and multiply to equal 144 and 18?

Answer by mathsmiles(68)

I'm going to guess you mean they multiply to 144 and they add to 18.

Let's choose X and Y for these two numbers.
X + Y = 18
X x Y = 144

Let's use substitution, which means we have to get either X or Y alone on one side of the equation:
X + Y = 18
X = 18 - Y

Now substitute this in place of X in the 2nd equation:
X x Y = 144
(18-y) x Y = 144 multiplying out the parens
18Y - Y^2 = 144 reordering:
-y^2 + 18Y = 144 subtracting 144 from both sides:
-y^2 + 18Y - 144 = 0

Just because I prefer to make life easier when possible, I'd like to get rid of the negative on the Y^2 term, so multiply both sides by -1.
-1(-y^2 + 18Y - 144) = -1(0)
y^2 - 18Y + 144 = 0
(Y - _)(Y - _) = 0
We need two factors of 144 whose sum is 18. Hmm...

Since two factors of 144 are 12 and 12, and other factors will be equally large, I don't see any way we're going to find two factors of 144 that add up to 18. This proves, I don't understand the problem the way you wrote it. Please correct it and re-submit.

All the details do matter. :-)